The 364, 308 was removed for bad behavior (jumps) just after TS5 mounting.

Tested gains of TS4: ok. (not cause shell to oscillate)

Ghosts analisys done.: towerdata/adsec_calib/noise/20100818_111254/noise_spectrum_ts4_gains.sav

Found 360-367 board with bad spectrum even with only low P gain.

full spectrum: /towerdata/adsec_calib/noise/20100818_111442/noise_spectrum_ts4_gains_full.sav

Step responses data of a small subset:

not working: there is some problem in the lower left side : found in the board 360-367: commands of 0.8 mm (!) (TN: 20100818_112318)

removed 360-367 from cl_act: step repetition: ok! (TN 20100818_123648)

Another FF acquired because the static error on the low order modes is very high.

Safe skip 2 times on mode 99: TN 20100818_124611, 20100818_125055

Done zonal ff, fitted and uploaded.

Start flattening: bug fixed in the trefoil trigger and increased the trigger amplitude to 12e-6 m

Slowly flattening for catch the ear

flattened 250 modes with 120nm rms residual (TTP removed) WAVEFRONT

Mask changed, fixing the RRH shadow.

Acquisition 0-250, 251-500 modes with 100e-9 m, reduced.

Flat 500 modes: WFE 124 nm (TTP removed):, CONVECTION RMS PEAK 60nm, mean 40nm

aggiungi immagini!!!

-- MarcoXompero - 18 Aug 2010
Topic revision: r4 - 19 Aug 2010, MarcoXompero
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback